OpenStudy (anonymous):

The sum of two numbers is 51, but the two numbers have a difference of 5. What are all possible combinations?

OpenStudy (anonymous):

x + y = 51 x - y = 5 solve these for x and y

OpenStudy (anonymous):

adding 2 equations gives 2x = 56 so x = 28 thus y = 51 - 28 = 23 so ome of the numbers is 23 and the other is 28

OpenStudy (anonymous):

Are there any others b/c I found those already? I have two more possible combinations?

OpenStudy (anonymous):

oh - i cant see how there can be another combination